OpenStudy (anonymous):

Help please! :) The answer would be C right? Four ounces of a 20% salt solution is mixed with 6 ounces of a 15% salt solution. What is the percent of salt in the new solution? A. 17% B. 17.5% C. 35% D. 85%

OpenStudy (anonymous):

b

OpenStudy (mathstudent55):

Let's calculate it.

OpenStudy (anonymous):

A

OpenStudy (mathstudent55):

4 ounces of 20% salt solution contains how much salt?

OpenStudy (anonymous):

I meant a, sorry

OpenStudy (mathstudent55):

4 * 0.2 = 0.8 That means 0.8 ounces of salt.

OpenStudy (anonymous):

yeah

OpenStudy (mathstudent55):

6 ounces of 15% salt solution contain 6 * 0.15 = 0.9 ounces of salt.

OpenStudy (anonymous):

4 x 20% = .8 6 x 15% = .9 10 (total weight ) x 1.7

OpenStudy (mathstudent55):

Now add the two amounts of salt: 0.8 + 0.9 = 1.7 ounces of salt.

OpenStudy (mathstudent55):

Since you have 4 ounces of one solution and 6 ounces of the other solution, the total amount of solution is 4 + 6 = 10 ounces. 1.7 ounces of salt in 10 ounces of solution is 1.7/10 * 100 = 17%

OpenStudy (mathstudent55):

The answer is A. 17%

OpenStudy (anonymous):

Thanks! :)

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (anonymous):

what's a smart score? i'm new to open study

OpenStudy (mathstudent55):

Click on a person's name and look at the person's detailed profile.